OpenStudy (anonymous):

can someone help me integrate this? preferable with integration by parts 0.5xe^(0.5x^2)

OpenStudy (kainui):

I think it's probably better to use a substitution instead, should be much easier that way. Try \(u=.5x^2\) that should work.

OpenStudy (anonymous):

\[\int\limits_{?}^{?}0.5^{?}\times \epsilon ^{0,5}\]

OpenStudy (anonymous):

http://www.integral-calculator.com/

OpenStudy (anonymous):

yup. it should be using substitution method

OpenStudy (anonymous):

It won't work by parts at all...

OpenStudy (amistre64):

it works fine by parts since a polynomial derives to 0

OpenStudy (amistre64):

u = .5 dv = x e^(.5x^2) dx du = 0 dx v = e^(.5 x^2) int [u dv] = uv - int [v du] int [ .5* xe^(.5x^2)] dx = .5*e^(.5 x^2) - 0 int [e^(.5 x^2)] dx =0 int [ .5* xe^(.5x^2)] dx = .5*e^(.5 x^2)